# How Far are VLMs from Visual Spatial Intelligence? A Benchmark-Driven Perspective

### v1

@@ 1 Introduction
Figure 1: Performance of SOTA Models on 23 Visual Spatial Reasoning Tasks (left). The evaluation reveals that the models have significant room for improvement, especially in tasks requiring precise numerical estimation, perspective taking, temporal information processing, and, particularly, spatial imagination. See Table II and Table III for detailed results. Comparison of Visual Spatial Reasoning and General VQA (Upper-right). While general VQA tasks primarily focus on extracting semantic information from images, VSR necessitates a deeper capacity to model and reason about spatial relationships. Data Formats and Task Settings for Visual Spatial Reasoning (Bottom-right). The evaluation includes 3 input formats and 23 task settings, covering three levels: Basic Perception, Spatial Understanding, and Planning.
@@
Visual Spatial Reasoning represents a fundamental cognitive ability in humans [1, 2]. It enables us to derive rich spatial information from observing the world, which is crucial for navigating and interacting with real-world environments. This capacity is not only vital for humans but is also indispensable for advancing key AI domains such as embodied intelligence and autonomous driving. Consequently, achieving human-level visual Spatial Reasoning in machines has been a long-standing pursuit in the research community [3, 4, 5, 6, 7, 8, 9, 10]. However, this objective remains highly challenging, primarily due to the inherent complexity of representing and reasoning about three-dimensional (3D) space.
@@
In recent years, Vision-Language Models (VLMs) [11, 12, 13, 14, 15, 16, 17, 18, 19, 4, 20, 21, 22] attract significant attention due to the powerful visual understanding and reasoning capabilities, bringing widespread applications in various Visual Question Answering (VQA) tasks [23, 24, 25]. Concurrently, an emerging body of work [26, 27, 28, 29] begins to apply VLMs to Visual Spatial Reasoning (VSR). However, VSR presents a distinct and more formidable set of challenges compared to general-purpose VQA, which often focuses on semantic-level understanding, as illustrated in Fig 1. While general VQA targets object recognition or attribute identification, VSR demands intricate reasoning about complex spatial dynamics, such as the relative positions, orientations, distances, and object motion. This need for deep spatial awareness is particularly exacerbated in multi-view and video contexts, which not only amplify the reasoning complexity but also increase the model’s propensity for hallucinations. Furthermore, this difficulty extends to data curation. Building robust VSR datasets is substantially more demanding, requiring precise spatial annotations and complex scene analysis that vastly exceed the requirements of traditional VQA benchmarks.
@@
Meanwhile, research in VSR advances rapidly, with current efforts concentrating on two primary fronts. On one hand, the community continues to drive innovation in input modalities [27], model architectures [28], training strategies [30], and reasoning mechanisms [29]. On the other hand, there is a dedicated push [31, 32, 33, 34] to develop higher-quality and more diverse datasets that enable comprehensive evaluation of VSR capabilities.
@@ 2.1 Research Scope
In recent years, with the development of Vision-Language Models [3, 35, 4, 36, 37, 38, 39, 40, 41, 42, 19, 43, 44, 45, 46, 47, 48, 48, 5] (VLMs) and generative models [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66], the realization of spatial intelligence has become increasingly promising. Spatial intelligence manifests in various aspects. First, agents perceive the 3D world through sensor inputs, gaining an understanding of its basic properties, followed by the comprehension of spatial relationships and physical laws. Furthermore, agents can interact with their environment, such as performing spatial navigation and manipulating objects. Additionally, the ability to create and imagine entirely new worlds is also a crucial aspect of spatial intelligence. The scope of spatial intelligence is vast, covering a wide range of tasks and applications, such as spatial reasoning, embodied artificial intelligence [67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79], and video world models [80, 81, 82, 83, 84, 85, 86, 87, 88].
For agents, image/video input is a readily accessible and cost-effective form of data. For biological organisms, vision is also a vital pathway for spatial modeling. Therefore, this paper focuses on the application of VLMs in spatial reasoning, which involves interpreting spatial information from images, multi-view inputs, or videos, including basic perception, spatial relationship understanding, and planning. It should be noted that tasks involving point clouds [89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99] as input for multimodal spatial reasoning or pure-text-based [100, 101, 92, 102, 103] spatial reasoning are not within the scope of this research. Additionally, while generative models [104, 105, 106, 107, 108, 109] also reflect spatial intelligence, they differ significantly in their modeling mechanisms and spatial representation form VLMs, and thus are not part of this discussion. Our focus is primarily on the general ability of VLMs in understanding spatial relationships, specific applications like vision-language action models [110, 111, 112, 113, 114, 115, 116, 117, 118, 119] and vision-language navigation models are not the main focus of our study.
@@ 2.2 Related Work
Several studies have provided comprehensive reviews of the development and evaluation frameworks of Vision-Language Models (VLMs). For example,  [120] systematically organizes the technological evolution, evaluation benchmarks, and application scenarios of VLMs, highlighting the potential performance fluctuations when handling complex tasks. The study further suggests introducing new evaluation metrics, emphasizing the importance of visual localization and multimodal understanding.
The study [121] investigates the integration of LLMs with 3D spatial understanding, analyzing challenges in data representation, model architecture, and evaluation metrics. It also recommends improvements in these three areas. It should be noted that the study in [121] primarily focuses on using 3D representations as input, rather than image/video input. Zha et al[122] similarly explores the spatial reasoning capabilities of LLMs, categorizing tasks based on input modalities such as images, point clouds, and hybrid modalities. Additionally, some studies [123, 124, 125] have revealed shortcomings in the spatial reasoning abilities of VLMs, citing issues such as the lack of effective spatial attention mechanisms or the limited ability of existing attention mechanisms to align with object positions. Sapkot et al[126] reviews the development of Vision-Language Action Models (VLA) and points out challenges in real-time control, dataset bias, and system integration when applying these models in practice. Researchers in [127] discusses the application of VLMs in autonomous driving scenarios, noting that the reliability of VLMs remains insufficient in complex traffic environments. Guo et al[128] summarizes the technological evolution and applications of VLMs, highlighting challenges in data quality and complex tasks within the biomedical field.
The scope of this study is on the spatial reasoning abilities of VLMs, specifically targeting pure visual capabilities, where the inputs are single images, multi-view images, or videos. Our aim is to review current advancements in visual spatial reasoning and categorize spatial reasoning into three cognitive levels: basic perception, spatial understanding, and planning. We provide a detailed and systematic analysis of existing task formulations. Additionally, we have gathered and curated a high-quality, comprehensive dataset from open-source benchmarks to evaluate the spatial reasoning abilities of VLMs.
@@ 3.1 Naive Solution
Generally, the vision encoder processes the image to extract a set of feature vectors $F_{v}=f_{\mathrm{vision}}(I)$, where $F_{v}\in\mathbb{R}^{N\times D_{v}}$ is a sequence of N feature vectors, each with dimension $D_{v}$, capturing both semantic and spatial information. Concurrently, the text question $Q$ is converted into embeddings $E_{q}\in\mathbb{R}^{M\times D_{t}}$ by text encoder. Given the potential mismatch in dimensions and feature spaces between the two modalities, the projection module is employed to align the visual features with the text embedding space. This module transforms $F_{v}$ into a new representation $F^{\prime}_{v}=g_{\mathrm{proj}}(F_{v})$. This projection is usually implemented with a few linear layers or a cross-attention mechanism, mapping the visual features so that $F^{\prime}_{v}\in\mathbb{R}^{N\times D_{t}}$, aligning their dimension with the text embeddings. Finally, the features from both modalities are concatenated and fed into the LLM, which then autoregressively generates the answer: $A=f_{\mathrm{LLM}}(\text{concat}(F^{\prime}_{v},E_{q}))$. Notable examples of this structure include SpatialVLM[129], LLaVA-VSD[17], and other related works[130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142]. This naive solution is both concise and efficient. Subsequent work, focusing on the specific characteristics of visual-spatial reasoning tasks, typically introduces improvements in four key aspects: input modalities, model architecture, training strategies, and inference methods.
@@ 3.2 Assisted Input Modalities
RGB images are a planar projection of the 3D world, providing continuous texture information. However, during the projection process, real 3D structural information is lost, which is one of the challenges in visual spatial reasoning tasks [143]. 3D point clouds offer a better representation of structure, but their data scale, diversity, and quality fall far behind that of images. As a result, some approaches [28, 27] attempt to strike a balance between 3D and 2D. These methods, often referred to as 2.5D [7], utilize depth maps as an additional modality. The motivation is that combining depth maps with images can yield a representation similar to point clouds, enabling the recovery of 3D structure in the real world, while the input images retain complete complex texture. Furthermore, such RGB-D data is relatively easy to obtain, especially in embodied scenarios.
@@
Due to the substantial distribution gap between depth maps and natural images, effectively encoding depth maps is by no means a trivial task. Additionally, the visual encoders of most VLMs are only trained on text-image pairs, and simply concatenating RGB and depth features may negatively affect performance [27].
@@
To address these issues, SpatialRGPT[27] first replicates the single-channel depth map $D$ to match the three channels of an RGB image. It then employs a shared vision encoder, $f_{\mathrm{vision}}$ to process both modalities, yielding features $F_{v}=f_{\mathrm{vision}}(I)$ and $F_{d}=f_{\mathrm{vision}}(D)$. These features are subsequently passed to different projectors, a design which ensures the model can still function normally without depth input and avoids the need for large-scale image-depth data training. Building on this, RoboRefer [28] uses a completely separate Depth encoder $f_{\mathrm{depth}}$ to handle depth inputs, thereby avoiding modality interference. This architecture choice can be expressed as: $F_{v}=f_{\mathrm{vision}}(I)$ and $F_{d}=f_{\mathrm{depth}}(D)$. This separation helps preserve general VQA performance while enhancing the perception of depth cues. SSR [144] also aims to integrate depth information in a similar manner. In addition to using modular designs for feature fusion, SpatialCLIP[143] is inspired by point cloud reconstruction and 3D point cloud architectures. It utilizes depth information to lift 2D tokens to 3D space, thereby better modeling the 3D spatial relationships. Specifically, for each transformer layer, it first lifts the 2D tokens to 3D voxels based on the average depth of each patch, and then introduces 3D depth convolutions to capture spatial patterns in the 3D space, thus providing enhanced spatial information.
@@
In addition to processing at feature space level, some models introduce an explicit, symbolic reasoning step by interacting with external tools. SpatialBot [145] provides a method for querying depth values from a given depth map $D$ through an API. The process begins when the model receives an image $I$ and a question $Q$. Based on the question, SpatialBot first determines if depth information is required. If so, it identifies the target coordinates $(x,y)$ in the image relevant to the question. For instance, when asked about an object’s depth, it first computes the objects’s bounding box, $B_{\mathrm{obj}}$, and then determines its center point: $(x_{c},y_{c})=\text{Center}(B_{\mathrm{obj}})$. Subsequently, the model generates a symbolic API call, such as $\text{Depth}(x_{c},y_{c})$. This triggers an external function that queries the depth map $D$ to retrieve the corresponding depth value $d$: $d=\text{API\_Query}(D,(x_{c},y_{c}))$. This numerical value $d$ is then converted back into a natural language string, $T_{d}$. Finally, SpatialBot generates the final answer A by processing a new augmented prompt that concatenates the original question with the retrieved depth information. This method of querying an API and re-inputting the information as text is further explored by MM-Spatial [146]. Their experiments also revealed that this “Query-and-Re-input” cycle yields significantly better results for spatial reasoning tasks in VLMs. While this class of methods often achieves high accuracy, its pipeline is also quite complicated.
@@ 3.3 Model Architecture Optimization
Numerous existing Vision-Language Models, such as LLaVA[147] and its subsequent versions [18, 140, 148] employ contrastive learning for pretraining visual encoders, generating compact and expressive visual embeddings. These models align well with natural language, demonstrating strong performance in tasks like image captioning and general visual question answering. However, similar to the contrastive learning used in CLIP[47], they primarily optimize for global semantic alignment and often neglect fine-grained spatial reasoning. Consequently, the visual embeddings produced by these encoders, while capturing an image’s overall semantic gist, provide only a coarse representation of its contents and are inherently limited in encoding precise spatial information.
To overcome this limitation, a powerful strategy has emerged: the dual-visual-encoder architecture. This design addresses the issue by creating two complementary visual pathways. While the original encoder provides the high-level semantic context, a second, specialized encoder is introduced to supply the fine-grained details. For single-image inputs, models like SpatialLLM[149] and ViCA2[150] incorporate additional encoders pre-trained on detail-oriented tasks. The reason for choosing models like MAE[151], DINO v2[152], and SAM[153] is that their training objectives, such as masked image reconstruction or large-scale segmentation, force them to learn rich, pixel-level features that the primary semantic encoder ignores. For inputs with inherent 3D information, like video or multi-view images, this approach is taken a step further. Standard fine-grained encoders may not be sufficient to interpret the geometric cues embedded in motion and parallax. Therefore, several studies use 3D reconstruction models as the second encoder. For instance, VLM-3R [154] incorporates CUT3R [155] to model the scene’s underlying 3D geometry from multiple views. By doing so, the model gains access to much richer representation that include depth and structure. Similarly, VG-LLM [156] and Spatial-MLLM[157] use advanced reconstruction networks like VGGT[158] to provide the language model with an even more sophisticated understanding of 3D space, significantly enhancing its spatial reasoning capabilities.

### v2

@@ 1 Introduction
Figure 1: Performance of SOTA Models on 23 Visual Spatial Reasoning Tasks (left). The evaluation reveals that the models have significant room for improvement, especially in tasks requiring precise numerical estimation, perspective taking, temporal information processing, and, particularly, spatial imagination. See Table I and Table II for detailed results. Comparison of Visual Spatial Reasoning and General VQA (Upper-right). While general VQA tasks primarily focus on extracting semantic information from images, VSR necessitates a deeper capacity to model and reason about spatial relationships. Data Formats and Task Settings for Visual Spatial Reasoning (Bottom-right). The evaluation includes 3 input formats and 23 task settings, covering three levels: Basic Perception, Spatial Understanding, and Planning.
@@
Visual Spatial Reasoning represents a fundamental cognitive ability in humans [piaget2013child, moser2015place]. It enables us to derive rich spatial information from observing the world, which is crucial for navigating and interacting with real-world environments. This capacity is not only vital for humans but is also indispensable for advancing key AI domains such as embodied intelligence and autonomous driving. Consequently, achieving human-level visual Spatial Reasoning in machines has been a long-standing pursuit in the research community [alayrac2022flamingo, dai2023instructblip, liu2025ola, chen2024end, jin2023adapt, bojarski2016end, kendall2019learning, chitta2022transfuser]. However, this objective remains highly challenging, primarily due to the inherent complexity of representing and reasoning about three-dimensional (3D) space.
@@
In recent years, Vision-Language Models (VLMs) [InternVL3, BAGEL, Qwen3, BLIP, Flamingo, Seed1.5-vl, jin2024llava, li2024llava, chen2025blip3, dai2023instructblip, gpt5, team2023gemini, zeng2025glm] attract significant attention due to the powerful visual understanding and reasoning capabilities, bringing widespread applications in various Visual Question Answering (VQA) tasks [kuang2025natural, albalak2025bigmathlargescalehighqualitymath, pennino2025reasoning]. Concurrently, an emerging body of work [SpatialVLM, SpatialRGPT, zhou2025roborefer, Visual-Cot] begins to apply VLMs to Visual Spatial Reasoning (VSR). However, VSR presents a distinct and more formidable set of challenges compared to general-purpose VQA, which often focuses on semantic-level understanding, as illustrated in Fig 1. While general VQA targets object recognition or attribute identification, VSR demands intricate reasoning about complex spatial dynamics, such as the relative positions, orientations, distances, and object motion. This need for deep spatial awareness is particularly exacerbated in multi-view and video contexts, which not only amplify the reasoning complexity but also increase the model’s propensity for hallucinations. Furthermore, this difficulty extends to data curation. Building robust VSR datasets is substantially more demanding, requiring precise spatial annotations and complex scene analysis that vastly exceed the requirements of traditional VQA benchmarks.
@@
Meanwhile, research in VSR advances rapidly, with current efforts concentrating on two primary fronts. On one hand, the community continues to drive innovation in input modalities [SpatialRGPT], model architectures [zhou2025roborefer], training strategies [ouyang2025spacer], and reasoning mechanisms [Visual-Cot]. On the other hand, there is a dedicated push [VSI-Bench, Spatial-MM, OmniSpatial, SpatialBench] to develop higher-quality and more diverse datasets that enable comprehensive evaluation of VSR capabilities.
@@ 2.1 Research Scope
In recent years, with the development of Vision-Language Models [alayrac2022flamingo, yang2023dawn, dai2023instructblip, chen2025blip3, wu2025omnigen2, team2025kwai, liu2025ola] (VLMs) and generative models [labs2025flux1kontextflowmatching, li2024hunyuan, dhariwal2021diffusion, song2020denoising, ho2020denoising, katzir2023noise], the realization of spatial intelligence has become increasingly promising. Spatial intelligence manifests in various aspects. First, agents perceive the 3D world through sensor inputs, gaining an understanding of its basic properties, followed by the comprehension of spatial relationships and physical laws. Furthermore, agents can interact with their environment, such as performing spatial navigation and manipulating objects. Additionally, the ability to create and imagine entirely new worlds is also a crucial aspect of spatial intelligence. The scope of spatial intelligence is vast, covering a wide range of tasks and applications, such as spatial reasoning, visual-language localization [chu2024towards, ju2024video2bev], embodied artificial intelligence [zitkovich2023rt, black2024pi_0, kim2024openvla, chi2024universal], and video world models [gao2024cardreamer, yang2025driving, zheng2024occworld].
For agents, image/video input is a readily accessible and cost-effective form of data. For biological organisms, vision is also a vital pathway for spatial modeling. Therefore, this paper focuses on the application of VLMs in spatial reasoning, which involves interpreting spatial information from images, multi-view inputs, or videos, including basic perception, spatial relationship understanding, and planning. It should be noted that tasks involving point clouds [pan2025omnimanip, ding2024holistic, li2024know, rizvi2024sparc, zhou2025physvlm, yang20253d, qi2024shapellm] as input for multimodal spatial reasoning or pure-text-based [li2024advancing, mouselinos2024beyond, rizvi2024sparc, wang2024dspy, li2025imagine] spatial reasoning are not within the scope of this research. Additionally, while generative models [bogdoll2023muvo, cai2023diffdreamer, yang2024worldgpt, bruce2024genie, liu2024world, yin2023nuwa] also reflect spatial intelligence, they differ significantly in their modeling mechanisms and spatial representation form VLMs, and thus are not part of this discussion. Our focus is primarily on the general ability of VLMs in understanding spatial relationships, specific applications like vision-language action models [liao2025diffusiondrive, li2025generative, jiang2025transdiffuser, tang2025hip] and vision-language navigation models are not the main focus of our study.
@@ 2.2 Related Work
Several studies have provided comprehensive reviews of the development and evaluation frameworks of Vision-Language Models (VLMs). For example,  [li2025benchmark] systematically organizes the technological evolution, evaluation benchmarks, and application scenarios of VLMs, highlighting the potential performance fluctuations when handling complex tasks. The study further suggests introducing new evaluation metrics, emphasizing the importance of visual localization and multimodal understanding.
The study [ma2024llms] investigates the integration of LLMs with 3D spatial understanding, analyzing challenges in data representation, model architecture, and evaluation metrics. It also recommends improvements in these three areas. It should be noted that the study in [ma2024llms] primarily focuses on using 3D representations as input, rather than image/video input. Zha et al[zha2025enable] similarly explores the spatial reasoning capabilities of LLMs, categorizing tasks based on input modalities such as images, point clouds, and hybrid modalities. Additionally, some studies [qi2025beyond, chen2025spatial, zhang2024vision] have revealed shortcomings in the spatial reasoning abilities of VLMs, citing issues such as the lack of effective spatial attention mechanisms or the limited ability of existing attention mechanisms to align with object positions. Sapkot et al[sapkota2025vision] reviews the development of Vision-Language Action Models (VLA) and points out challenges in real-time control, dataset bias, and system integration when applying these models in practice. Researchers in [xie2025vlms] discusses the application of VLMs in autonomous driving scenarios, noting that the reliability of VLMs remains insufficient in complex traffic environments. Guo et al[guo2024survey] summarizes the technological evolution and applications of VLMs, highlighting challenges in data quality and complex tasks within the biomedical field.
@@ 3.1 Naive Solution
Generally, the vision encoder processes the image to extract a set of feature vectors $F_{v}=f_{\mathrm{vision}}(I)$, where $F_{v}\in\mathbb{R}^{N\times D_{v}}$ is a sequence of N feature vectors, each with dimension $D_{v}$, capturing both semantic and spatial information. Concurrently, the text question $Q$ is converted into embeddings $E_{q}\in\mathbb{R}^{M\times D_{t}}$ by text encoder. Given the potential mismatch in dimensions and feature spaces between the two modalities, the projection module is employed to align the visual features with the text embedding space. This module transforms $F_{v}$ into a new representation $F^{\prime}_{v}=g_{\mathrm{proj}}(F_{v})$. This projection is usually implemented with a few linear layers or a cross-attention mechanism, mapping the visual features so that $F^{\prime}_{v}\in\mathbb{R}^{N\times D_{t}}$, aligning their dimension with the text embeddings. Finally, the features from both modalities are concatenated and fed into the LLM, which then autoregressively generates the answer: $A=f_{\mathrm{LLM}}(\text{concat}(F^{\prime}_{v},E_{q}))$. Notable examples of this structure include SpatialVLM[chen2024spatialvlm], LLaVA-VSD[jin2024llava], and other related works[liu2025coarse, wen2024can, tang2024sparkle, liu2022things, ogezi2025spare, wang2024root, yuan2024robopoint, liu2025reasongrounder, team2025robobrain, xu2025multi, xu2025llava, li2024proximity, zhao2024can]. This naive solution is both concise and efficient. Subsequent work, focusing on the specific characteristics of visual-spatial reasoning tasks, typically introduces improvements in four key aspects: input modalities, model architecture, training strategies, and inference methods.
@@ 3.2 Assisted Input Modalities
RGB images are a planar projection of the 3D world, providing continuous texture information. However, during the projection process, real 3D structural information is lost, which is one of the challenges in visual spatial reasoning tasks [wang2025spatialclip]. 3D point clouds offer a better representation of structure, but their data scale, diversity, and quality fall far behind that of images. As a result, some approaches [zhou2025roborefer, SpatialRGPT] attempt to strike a balance between 3D and 2D. These methods, often referred to as 2.5D [jin2023adapt], utilize depth maps as an additional modality. The motivation is that combining depth maps with images can yield a representation similar to point clouds, enabling the recovery of 3D structure in the real world, while the input images retain complete complex texture. Furthermore, such RGB-D data is relatively easy to obtain, especially in embodied scenarios.
@@
Due to the substantial distribution gap between depth maps and natural images, effectively encoding depth maps is by no means a trivial task. Additionally, the visual encoders of most VLMs are only trained on text-image pairs, and simply concatenating RGB and depth features may negatively affect performance [SpatialRGPT].
@@
To address these issues, SpatialRGPT[SpatialRGPT] first replicates the single-channel depth map $D$ to match the three channels of an RGB image. It then employs a shared vision encoder, $f_{\mathrm{vision}}$ to process both modalities, yielding features $F_{v}=f_{\mathrm{vision}}(I)$ and $F_{d}=f_{\mathrm{vision}}(D)$. These features are subsequently passed to different projectors, a design which ensures the model can still function normally without depth input and avoids the need for large-scale image-depth data training. Building on this, RoboRefer [zhou2025roborefer] uses a completely separate Depth encoder $f_{\mathrm{depth}}$ to handle depth inputs, thereby avoiding modality interference. This architecture choice can be expressed as: $F_{v}=f_{\mathrm{vision}}(I)$ and $F_{d}=f_{\mathrm{depth}}(D)$. This separation helps preserve general VQA performance while enhancing the perception of depth cues. SSR [liu2025ssr] also aims to integrate depth information in a similar manner. In addition to using modular designs for feature fusion, SpatialCLIP[wang2025spatialclip] is inspired by point cloud reconstruction and 3D point cloud architectures. It utilizes depth information to lift 2D tokens to 3D space, thereby better modeling the 3D spatial relationships. Specifically, for each transformer layer, it first lifts the 2D tokens to 3D voxels based on the average depth of each patch, and then introduces 3D depth convolutions to capture spatial patterns in the 3D space, thus providing enhanced spatial information.
@@
In addition to processing at feature space level, some models introduce an explicit, symbolic reasoning step by interacting with external tools. SpatialBot [cai2024spatialbot] provides a method for querying depth values from a given depth map $D$ through an API. The process begins when the model receives an image $I$ and a question $Q$. Based on the question, SpatialBot first determines if depth information is required. If so, it identifies the target coordinates $(x,y)$ in the image relevant to the question. For instance, when asked about an object’s depth, it first computes the objects’s bounding box, $B_{\mathrm{obj}}$, and then determines its center point: $(x_{c},y_{c})=\text{Center}(B_{\mathrm{obj}})$. Subsequently, the model generates a symbolic API call, such as $\text{Depth}(x_{c},y_{c})$. This triggers an external function that queries the depth map $D$ to retrieve the corresponding depth value $d$: $d=\text{API\_Query}(D,(x_{c},y_{c}))$. This numerical value $d$ is then converted back into a natural language string, $T_{d}$. Finally, SpatialBot generates the final answer A by processing a new augmented prompt that concatenates the original question with the retrieved depth information. This method of querying an API and re-inputting the information as text is further explored by MM-Spatial [daxberger2025mm]. Their experiments also revealed that this “Query-and-Re-input” cycle yields significantly better results for spatial reasoning tasks in VLMs. While this class of methods often achieves high accuracy, its pipeline is also quite complicated.
@@ 3.3 Model Architecture Optimization
Numerous existing Vision-Language Models, such as LLaVA[liu2023visual] and its subsequent versions [li2024llava, xu2025llava, guo2024llava] employ contrastive learning for pretraining visual encoders, generating compact and expressive visual embeddings. These models align well with natural language, demonstrating strong performance in tasks like image captioning and general visual question answering. However, similar to the contrastive learning used in CLIP[radford2021learning], they primarily optimize for global semantic alignment and often neglect fine-grained spatial reasoning. Consequently, the visual embeddings produced by these encoders, while capturing an image’s overall semantic gist, provide only a coarse representation of its contents and are inherently limited in encoding precise spatial information.
To overcome this limitation, a powerful strategy has emerged: the dual-visual-encoder architecture. This design addresses the issue by creating two complementary visual pathways. While the original encoder provides the high-level semantic context, a second, specialized encoder is introduced to supply the fine-grained details. For single-image inputs, models like SpatialLLM[ma2025spatialllm] and ViCA2[feng2025towards] incorporate additional encoders pre-trained on detail-oriented tasks. The reason for choosing models like MAE[he2022masked], DINO v2[oquab2023dinov2], and SAM[kirillov2023segment] is that their training objectives, such as masked image reconstruction or large-scale segmentation, force them to learn rich, pixel-level features that the primary semantic encoder ignores. For inputs with inherent 3D information, like video or multi-view images, this approach is taken a step further. Standard fine-grained encoders may not be sufficient to interpret the geometric cues embedded in motion and parallax. Therefore, several studies use 3D reconstruction models as the second encoder. For instance, VLM-3R [VSTiBench] incorporates CUT3R [wang2025continuous] to model the scene’s underlying 3D geometry from multiple views. By doing so, the model gains access to much richer representation that include depth and structure. Similarly, VG-LLM [zheng2025learning] and Spatial-MLLM[wu2025spatial] use advanced reconstruction networks like VGGT[wang2025vggt] to provide the language model with an even more sophisticated understanding of 3D space, significantly enhancing its spatial reasoning capabilities.